# ProEx: A Unified Framework Leveraging Large Language Model with Profile Extrapolation for Recommendation

### v1

@@ A.2. Multi-faceted Profile Construction
A.3. Experimental Details
A.3.1. Datasets
Detailed statistics for three datasets are presented in Table 4. Each dataset is divided into training, validation, and test sets following a 3:1:1 split ratio. The last column shows the number of user and item profiles for each dataset, where each user (item) has one original profile (OP) and three new profiles (NP #1, NP #2, and NP #3).
A.3.2. Base Model Introduction
@@ A.3.3. LLM-enhanced Baselines
Table 4. Statistics of the datasets.
Dataset #Users #Items #Interactions #Profiles
Amazon-Book 11,000 9,332 200,860 81,328
Yelp 11,091 11,010 277,535 88,404
Steam 23,310 5,237 525,922 114,188
A.3.4. Detailed Hyperparameter Settings
@@
For ProEx, the number of profiles $\mathcal{K}$ is fixed to 4, including one original profile and three newly generated ones (see Fig. 9 for details). The number of environments $\mathcal{E}$ ranges from 2 to 8 (see Fig. 8 for details). The shape parameter $\alpha_{k}$ required for sampling in the linear combination within each environment (as defined in Eq. 10) is selected from [0.1,0.2,0.3], with 0.1 yielding the best performance in most cases. The temperature coefficient $\tau$ in the contrastive regularization loss $\mathcal{L}_{\text{reg}}$ (Eq. 7) is set to 0.2 by default. For the cross-space alignment loss, we use contrastive loss for discriminative methods, with the temperature coefficient set to 0.2 by default. For generative methods, we adopt the distribution matching approach defined in Eq. 11, where the value of $\beta$ is selected from [0.0, 1.0] with a step size of 0.1. The weight $\lambda_{1}$ of the alignment loss is selected from the range [0,1], while the weight of the contrastive regularization loss $\lambda_{2}$ is chosen from [0,2] with a step size of 0.1. Empirically, we find that setting the weight of the variance term $\lambda_{3}$ in Eq. 14 to 1.0 yields good performance in most cases.
LightGCN Mult-VAE
R@20 N@20 R@20 N@20
base 0.1411 0.0856 0.1472 0.0916
(a) Amazon-Book
(b) Yelp
Figure 11. Comparison of recommendation performance between ProEx and DMRec based on different profiles w.r.t. Recall@20 and NDCG@20 on (a) Amazon-Book dataset and (b) Yelp dataset.
(a) Yelp
(b) Steam
Figure 12. Comparison of the base model and ProEx performance across sparse user groups. The bar chart (left $y$-axis) shows the number of users in each sparse user group, and the line chart (right $y$-axis) shows NDCG@20.
A.4. Supplementary Experiments
A.4.1. Impact of Different Semantic Representations
Transforming the generated profiles into representations in the feature space is a crucial step in ProEx, which is primarily accomplished through the text embedding model $f_{\text{text}}$. To investigate the impact of different text embedding models on ProEx, we conduct experiments using various semantic representations on the base models LightGCN (he2020lightgcn) and Mult-VAE (liang2018variational), as shown in Table 5.
Firstly, all types of semantic representations improve the performance of the base models to varying degrees, further confirming the effectiveness and generalizability of ProEx. Secondly, despite LLama3 generating the largest output dimension (4096), it yields the poorest performance, especially in the case of LightGCN. This may be attributed to the limited capability of an 8B-scale LLM in comprehending input text, thereby producing low-quality semantic representations. In addition, the substantial dimensionality discrepancy (e.g., the embedding size of LightGCN is merely 32) exacerbates the challenge of dimension alignment. In contrast, the locally deployable SFR-Embedding-Mistral achieves surprisingly strong results, indicating that ProEx can attain good recommendation performance without necessarily relying on closed-source LLMs (e.g., GPT).
A.4.2. Extended Comparison against DMRec
Recalling Fig. 4(b), we present the performance of the original profile and the three new profiles in DMRec (zhang2025towards), which shows significant fluctuations. In this section, we further compare the performance of DMRec and ProEx on the Amazon-Book and Yelp datasets. For DMRec, we conduct tests using different profiles, and the results are shown in Fig. 11. It can be observed that DMRec still exhibits noticeable performance fluctuations on the Yelp dataset, which strongly supports the two issues mentioned in the Introduction section: (1) the flaws of a single profile may adversely affect downstream recommendation tasks; and (2) it is difficult to intuitively assess the quality of each profile. In contrast, ProEx can fully leverage the information from multiple profiles to collaboratively enhance recommendation performance, thereby reducing reliance on the quality of any single profile.
To examine whether ProEx can alleviate the data sparsity issue, we perform a sparsity analysis on Mult-VAE, DMRec, and ProEx. Following the approach in (zhang2023revisiting; yu2022graph), users are divided into four groups {U1, U2, U3, U4} based on their interaction scales, and each group is evaluated separately. The results of these experiments are presented in Fig. 12. Let $s(\text{U})$ be the sparsity level of the user group U, then we have $s(\text{U1})>s(\text{U2})>s(\text{U3})>s(\text{U4})$ Intuitively, the proportion of users in the sparsest group is remarkably high (over 50% in both datasets), highlighting the extreme sparsity of these recommendation scenarios. Both DMRec and ProEx achieve performance improvements in the sparsest group, demonstrating the effectiveness of incorporating LLMs to alleviate recommendation sparsity. Building on this, ProEx further improves performance, primarily benefiting from the collaborative effect of multiple profiles.
(a) LightGCN
(b) Mult-VAE
Figure 13. Impact of the weight of regularization loss $\lambda_{2}$ for (a) LightGCN and (b) Mult-VAE w.r.t. Recall@20 and NDCG@20 on Amazon-Book dataset.
A.4.3. Impact of the Weight of Regularization Loss $\lambda_{2}$
Finally, we investigate the weight of the contrastive regularization loss $\lambda_{2}$ used to enhance profile extrapolation. Similarly, we present the variation of $\lambda_{2}$ on Amazon-Book dataset when using LightGCN (he2020lightgcn) and Mult-VAE (liang2018variational) as base models, as shown in Fig. 13. It can be observed that Mult-VAE generally requires a larger weight, whereas LightGCN only needs to be examined within a smaller range. This observation is reasonable, as LightGCN employs a pairwise recommendation loss (rendle2009bpr) that is typically small (less than 1), where an overly large regularization term could disrupt the training process. Conversely, Mult-VAE relies on a reconstruction error that is generally much larger (ranging from tens to hundreds), thereby necessitating a higher weight to balance the two losses on a comparable scale.

### v2

@@ A.1. Multi-faceted Profile Construction
A.2. Experimental Details
A.2.1. Base Model Introduction
@@ A.2.2. LLM-enhanced Baselines
A.2.3. Detailed Hyperparameter Settings
@@
For ProEx, the number of profiles $K$ is fixed to 4, including one original profile and three newly generated ones (see Fig. 9 for details). The number of environments $\mathcal{E}$ ranges from 2 to 8 (see Fig. 8 for details). The shape parameter $\alpha_{k}$ required for sampling in the linear combination within each environment (as defined in Eq. 10) is selected from [0.1,0.2,0.3], with 0.1 yielding the best performance in most cases. The temperature coefficient $\tau$ in the contrastive regularization loss $\mathcal{L}_{\text{reg}}$ (Eq. 7) is set to 0.2 by default. For the cross-space alignment loss, we use contrastive loss for discriminative methods, with the temperature coefficient set to 0.2 by default. For generative methods, we adopt the distribution matching approach defined in Eq. 11, where the value of $\beta$ is selected from [0.0, 1.0] with a step size of 0.1. The weight $\lambda_{1}$ of the alignment loss is selected from the range [0,1], while the weight of the contrastive regularization loss $\lambda_{2}$ is chosen from [0,2] with a step size of 0.1. Empirically, we find that setting the weight of the variance term $\lambda_{3}$ in Eq. 14 to 1.0 yields good performance in most cases.